OpenStudy (anonymous):

Which statement is equivalent to |6x + 2| ≤ 45?

OpenStudy (anonymous):

6x + 2 ≤ 45 6x + 2 ≤ -45 -45 ≤ 6x + 2 ≤ 45 6x + 2 ≤ 45 or 6x + 2 ≥ -45

OpenStudy (jackellyn):

|6x+2| <= 45 is the same as 6x+2 <= 45 and 6x+2<= -45

OpenStudy (jackellyn):

I would say the third one.